OpenStudy (anonymous):

PLEASE HELP SIGMA NOTATION Evaluate the summation of negative 5 n minus 1, from n equals 3 to 12.. −410 −385 385 410

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Compassionate

OpenStudy (freckles):

so you are going to add terms that look like 5n-1 starting at n=3 all the way to n=12

OpenStudy (freckles):

or if you don't like that you could use \[\sum_{n=1}^{k}n=\frac{k(k+1)}{2} \\ \sum_{n=1}^{k}c=ck\] if you rewrite your sum as: \[\sum_{n=1}^{12}(5n-1) - [5(1)-1]-[5(2)-1]\]

OpenStudy (freckles):

either way